# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

### v1

@@ 2 Literature review
Due to the recent introduction of Energy Communities, few papers address the specific mechanisms by which RECs interact with the wholesale market. Few theoretical studies based on narratives (Robinson and Arcos-Vargas, 2023; Robinson and Del Guayo, 2022; Williamson, 2022) speculate briefly on the direction of the systemic impact of RECs on wholesale quantity and prices. Moreover, we have found only four empirical studies on the matter (Backe et al., 2022; Fuentes González et al., 2022; Sarfarazi et al., 2020; Boccard and Goetz, 2025). Several empirical studies quantify the effect of individual prosumers (e.g.(Chen et al., 2023a, b; Tsybina et al., 2023)) as well as individual and aggregated prosumers (Riaz et al., 2019) on wholesale power market operations. However, quantifying the effect of prosumers on the wholesale market is not the core objective of these studies. Another limitation of these works is that the authors test their models on a small subset of actual markets (usually only a limited number of nodes/buses) over a short time scale (e.g. one week).
The work of Riaz et al. (2019) on the effect of aggregated prosumers on the wholesale market is conceptually the closest to the approach adopted in the present paper. However, apart from the methodological approach, a critical conceptual difference is that the optimization model in Riaz et al. (2019) was tested on residential prosumers, whereas our mixed model uses outputs from the operation of various categories of prosumers united in a REC – including residential, public, non-profit organizations, and small and medium enterprises (SMEs) – as well as producers. Faia et al. (2021) take an opposite approach compared to our paper. The authors explore the effect of the wholesale market on prosumers’ bills by optimizing the minimum electricity volume required for prosumers to participate in both the wholesale and retail markets, suggesting that if households are aggregated (e.g. by a REC), they can achieve significant cost reductions.
This Section reviews the relevant literature on the topic, by exploring two main areas:
@@
We have also reviewed studies that analyze the effect of prosumers on electricity systems and discuss their relevant findings in Section 5. However, since they do not constitute the central role in our paper, unlike studies that focus on the effect of REC deployment on the system, we omit reporting them in the main text although including them in the Appendix 7.3.
@@ 2.1 Impact of RECs on electricity systems
Several authors use a theoretical narrative-based approach to discuss the impacts of RECs (Biggar and Hesamzadeh, 2022; Del Pizzo et al., 2022; Parag and Sovacool, 2016; Robinson and Del Guayo, 2022; Di Silvestre et al., 2021). Robinson and del Guayo (2022) analyse REC features that potentially favour the system and its stakeholders, as well as those that may be harmful, using Spain as the case study. Biggar and Hesamzadeh (2022) emphasize the regulatory challenges arising from RECs and discuss various policy responses. Del Pizzo et al. (2022) examine criteria for REC sizing and geographical boundaries from a DSO’s perspective, aiming to align them with the needs of other grid users. Parag and Sovacool (2016) analyze the benefits and caveats of three categories of energy communities: peer-to-peer REC, microgrid REC, and aggregator REC. Di Silvestre et al. (2021) review REC regulations around the world and outline possible challenges and opportunities for power systems that arise from REC proliferation.
@@
Several studies indicate systemic benefits of REC deployment. Encouraging investment in embedded generation at a larger scale increases economic efficiency (Biggar and Hesamzadeh, 2022). Locating RECs in places with under-supply of electricity reduces transmission grid costs and losses due to an increased energy supply where it is mostly needed (Robinson and Del Guayo, 2022). RECs reduce the need for an additional large-scale generation due to the merit-order effect, which in turn reduces the wholesale price and the need for additional transmission infrastructure (Robinson and Del Guayo, 2022; Backe et al., 2022). Robinson and Guayo (2022) argue that RECs can provide ancillary services if aggregated into virtual power plants, thereby replacing conventional sources which are more expensive to operate in many cases. When RECs are aggregated and utilize BESS, they could reduce peak heat load and total heat demand in highly electrified systems (Backe et al., 2022). New ESCOs and aggregators could emerge as facilitators between RECs and the grid, therefore enhancing residential and commercial energy efficiency efforts (Parag and Sovacool, 2016). Finally, multiple social benefits (Biggar and Hesamzadeh, 2022; Robinson and Del Guayo, 2022) and improved competitive pressure on other market agents (Robinson and Del Guayo, 2022) are also important systemic benefits.
@@
At the same time, other studies point to systemic detriments of REC deployment. Biggar and Hesamzadeh (2022) and Robinson and Guayo (2022) argue that a REC may arbitrate [^9] Arbitrage effect emerges when collective generators and consumers/prosumers have distinct selling and buying prices. When they unite into an energy community, generator now can supply energy directly to members of the REC while latter remunerate the former with a payment that is higher than conventional price for selling into the grid but lower than a conventional buying price for consumers. As a result, both parties win while the regulatory effect of the price has vanished. between retail and selling prices in cases when it acts as a united entity which purchases part of its electricity externally from a single retailer. As a consequence, the effect of non-uniform tariffs (time-of-use and locational tariffs) is eliminated, which could make the system operation costlier and potentially may lead to the further expansion of the network. Biggar and Hesamzadeh (2022) indicate that retail customers do not typically face the correct incentives to use energy according to market and grid situations. When customers unite into RECs it may further amplify the impact of such inefficient tariffs. The same study (2022) argues that the reduction in network exploitation exacerbates network operator’s ‘death spiral’ effect [^10] Death spiral effect is a well-known phenomenon which describes a situation when distributed energy resources lead to prosumers paying less with volumetric distribution charges. Consequently, system operators (DSOs/TSOs) become under-paid for its previous infrastructure investment, which is usually higher than a decreased usage of a network due to self-consumption. System operators shift these costs instead onto other consumers, which in turn stimulate them to transform into prosumers as well, thereby shrinking the customer base iteratively.. Robinson and Guayo (2022) and Parag and Sovacool (2016) say that locating RECs in places with over-supply of electricity increases transmission grid costs and losses due to increased electricity injection where congestion bottlenecks are present. In addition, small capacity renewable generators are costlier for the system because they are more expensive by LCOE than large capacity renewable generators (Robinson and Del Guayo, 2022). Parag and Sovacool (2016) point that RECs might erode sensitive protections on privacy. Lastly, the same authors argue that some REC business models might be harder to deploy than others. For example, it is difficult to align the interests of peer-to-peer RECs with the interests of a broader society. Peer-to-peer business model can also entail high transaction costs. As we can see, the literature points to multiple potential harms to the system and stakeholders [^11] Since our study aims to analyze systemic impacts, we skip the literature review of impacts on the individual stakeholders in the main text. Instead, we include the review of these impacts in Appendix B. from the deployment of RECs. However, these potential harms relate to a variety of REC types and not all are necessarily applicable to the ‘community collective generation’ type deployed in Italy  [^12] Koltunov (2025a) can be consulted for an in-depth discussion on the impacts of individual REC business models on the electricity system and its stakeholders..
@@
Fuentes Gonzalez et al. (2022) utilize a game theoretical approach with optimization modelling for a simplified three-nodal Chilean market to find the potential impact of REC deployment on a consumer surplus, nodal prices, social welfare, generation and transmission expansion, and CO2 emissions. The study reveals that the nodal price decreases from 65.77 \$/MWh to 65.4 \$/MWh when RECs are deployed in a scenario with 9.24 MW of installed capacity with 150 MW peak nodal load. Moreover, the price drops further up to 64 \$/MWh if a REC’s installed capacity gets higher, though contingent upon the disposable income of REC members. The reduction in the nodal price suggests that the effect of RECs can also be positive on the non-member consumers. An increase in social welfare can be observed when RECs are involved in energy production (14,221 \$/hour with RECs versus 14,596 \$/hour without REC) due to the avoidance of high generation costs at nodes without REC, as well as the less frequent need for transmission expansion. This leads to a reduction in the nodal price that triggers an increase in the quantity demanded from 105.79 MWh to 115.03 MWh with an energy community. Quantities demanded in their static model are simulated for a single representative hour while prices are simply derived from linear demand functions. Consequently, this approach does not take into consideration a full interplay of technologies and variability over multiple time periods that are inherent in pay-as-clear zonal wholesale auctions of the real world.
@@
Sarfarazi et al. (2020) also use a game-theoretical approach with optimization and dynamic programming to simulate the interaction of a stylized German REC (that consists of flexible consumers, inflexible consumers, prosumers, prosumagers, and community energy storage) with its retailer and a wholesale market. The study indicates when community storage is owned by the retailer who has the profit maximization objective the scenarios with real-time purchase and feed-in selling tariffs do not decrease the quantities dispatched at the supply side of the wholesale market by RECs compared to the scenario with static tariffs when community storage decreases energy dispatched at the supply side of the wholesale market. In addition, in this scenario REC operations become aligned with the wholesale market price signals. Moreover, real-time tariffs with installed community energy storage increase profits for REC members. Therefore, this scenario leads to welfare gains both for the system and a REC. Conversely, if a retailer pursues a REC self-sufficiency maximization objective and not a profit maximization objective, REC imports less electricity from the public grid, therefore, less of grid charges are collected for the DSOs and less of levies to the general budget. Sarfarazi et al. (2020) focus primarily on the ”REC-retailer” interactions on a micro scale and do not investigate REC deployment on a macro scale, MOE and complexities of a DA wholesale market.
@@
Boccard and Goetz (2025) test empirically the profit sharing rules for a condominium energy community and a REC in three geographical locations with diverse solar irradiation levels [^13] Weak irradiation in Hamburg (Germany), medium irradiation in Girona (Northern Spain), high - in Faro (Southern Portugal) and for different PV sizes that correspond to 25%, 50%, and 100% load coverage of households in a REC. The study finds that REC members decrease their bills in all simulated scenarios. However, the impact on the grid can be diverse. When a PV system is sized at 25% of the aggregated load of REC members, power exchanges with the grid drop from 70 MWh (no PV scenario) to 59 MWh in Northern Spain, which is a 14% decrease compared to the scenario without a PV system. This is a significant finding which indicates that a DSO could manage the deployment of RECs and generally DERs without having to raise tariffs and without compromising on grid reliability. When the PV capacity is increased to cover 50% of the aggregated load, power exchanges with the grid decrease by 7%. In contrast, in the scenario where a PV capacity covers 100% of the member load, power exchanges with a distribution grid raise up to 29%. The situation worsens in the northern irradiance region (Germany), while improves in the southern irradiance location (Portugal). The opposite pattern was revealed by the sensitivity analysis. In the scenario with heterogeneous public buildings (schools and hospitals) that install PV systems covering 50% and 100% of the demand, the power exchanges reduce at 8% and 13%, respectively, compared to no PV scenario. Conversely, the 25% PV coverage of public loads raises power exchanges with the grid at 10%. A similar situation occurs with the higher number of more heterogeneous residential loads. The authors argue that the deployment of small-scale PV systems in residential RECs could decrease operation costs for the grid, while large-scale PVs force costly upgrades within buildings and across a distribution network. Therefore, the sizing of PV systems in RECs should be carefully calibrated not only from the perspective of economic benefits for REC members but also considering systemic impacts.
@@
Last but not least, Backe et al. (2022) utilized two large optimization models to explore the effect of the REC deployment on the additional capacity needs in a 2060 carbon-neutral Europe. The authors find that the REC deployment reduces both electricity and heating costs and lessens the need for national capacity expansion by 50-60 GW in six EU countries. However, they do not quantify the merit-order effect (MOE) or REC interaction with the DA market, openly stating that this is beyond the scope of their paper.
@@ 3 Methods and data
For clarity, we divided our methodological approach into two main stages. The first stage begins with mapping all Italian RECs to derive key input parameters for the subsequent engineering modeling. This model calculates both the energy injected into the grid and the energy self-consumed by various categories of prosumer load profiles, based on irradiation levels across different market zones. The model’s output is used to project REC deployment for the year 2024 (ex-post situation) and for the year 2027 (predictive situation), according to three main scenarios: best case, mid-case, and worst case. This projection step provides the input variables for the second stage of analysis.
In the second stage, we first simulate a market equilibrium with RECs that reflects the actual market situation, using Italian day-ahead market bids for the year 2024. Then, we construct a counterfactual scenario of market equilibrium without RECs for both 2024 and 2027. As the most recent annual data is available only for 2024, and the policy target is set for 2027, we assume the wholesale market conditions in 2027 to be identical to those of 2024 for simplification purposes. This assumption does not compromise our modeling objective, which is to introduce a new methodology for estimating the impact of REC deployment on wholesale market equilibrium. Our aim is not to forecast the actual impact of REC deployment in the policy-target year. Nonetheless, policy-relevant insights can still be drawn from the predicted situation. In the first stage of methodology, we used a combination of software applications: MS Excel (mapping), Matlab (engineering modeling), R (projection). In the second stage, R was utilized as the main modeling software. Each stage is described in more detail below.
@@ 3.1.1 Comprehensive Mapping of Italian RECs
  • the data portal of GSE [^15] The state-owned renewable energy agency
  • the annual reports of Legambiente [^16] A national environmental non-governmental organisation, i.e. NGO.
@@
  • the data portal NeXt [^17] NeXt — the civil society network comprising the majority of Italian third-sector and public bodies working in the REC field ESG
@@
  • the data portal Sinergie Condivise [^18] Developed by a banking foundation ‘Fondazione Compagnia di San Paolo‘ in collaboration with regional and municipal governments and universities in the Piedmont Region
  • publicly available business plans of individual RECs
  • academic publications
  • websites of REC developers
@@
  • websites of news agencies, regional and municipal authorities, and individual initiatives.
Our final database contains 34 variables for 362 RECs in Italy, of which 184 are in the operational phase and 178 are in the design phase as of 30 January 2025. However, for the purposes of this study, we used only the following 20 variables:
  • electricity market zone
  • number of primary substations to which the REC is connected
  • capacity (installed or planned), kW
  • PV capacity (installed or planned), kWp
  • capacity per substation, kW
  • self-consumption level, %
  • battery availability and capacity, kWh
  • number of members and installed capacity per category of prosumer/consumer/producer
  • building types per category of prosumer
  • capacity share of public prosumers in total REC capacity, %
  • capacity share of residential prosumers in total REC capacity, %
  • capacity share of SME prosumers in total REC capacity, %
  • capacity share of NPO prosumers in total REC capacity, %
  • capacity share of PV producers in total REC capacity, %
  • number of rooftop installations by public prosumers
  • number of rooftop installations by residential prosumers
  • number of rooftop installations by SME prosumers
  • number of rooftop installations by NPO prosumers
  • number of standalone PV installations
  • installed renewable energy technology.
Zhu et al. (2025) also built the database of new Italian RECs for the purpose of their study. In addition to the diverse scope of variables collected (due to different research objectives), another major difference lies in the development status of the RECs that were included in both databases. Our database includes both operational RECs and those that were in the project design stage of development as of January 2025. Inclusion of RECs in the project design stage allowed us to have a more holistic perspective on the future trends, therefore, to construct robust scenarios grounded in the larger sample of the real-world data, which is discussed in further detail in Section 3.1.3. The detailed comparison of two databases can be found in Table 3.1.1.
@@
Comparison between the databases Variable Our database Zhu et al. (2025) Total number of RECs, from which: 362 212 Operational RECs 184 212 Design-phase RECs 178 0 Location 362 212 Population (of municipality) 85 Climatic zone 85 Market zone 362 Generator’s capacity 300 212 Type of technology installed/planned 323 85 Number of membersa 283 Category of membersb 219 Share of installed PV capacity, by category of prosumers 271 Technical indicatorsc 147 BESS availability 13 7 EVs and/or e-charging availability 38 3 Self-consumption level 82 Economic indicatorsd 105 Number of prosumer buildings by type 158 55 Number of consumer buildings not counted 41 Investment source 252 CAPEX costs 106 Legal form 180 REC builders and promoters 204 Last update January 2025 February 2025 Data sources GSE, Legambiente, RSE, Sinergie, Condivise, NeXt ESG, business plans, academic publications, websites of builders, other websites GSE, Legambiente, RSE, academic publications
@@
From our database, we obtained input parameters for the subsequent methodological stages. The first input parameter can be calculated via the following equation:
$p_{a,one}^{P_{as}}=\frac{p_{a,total}^{P_{as}}}{N_{r}^{P_{as}}/N_{p}^{P_{as}}}$ (1)
@@
  • $p$ denotes a category of prosumer/producer.
  • $z$ denotes an electricity market zone.
  • $Z$ is the set of electricity market zones, where $Z\in\{\text{North},\text{Central North},\text{Central South},\text{Calabria},\\ \text{South},\text{Sicily},\text{Sardinia}\}$. [^19] From now, we refer to market zones through their codes NORD, CNORD, CSUD, CALA, SUD, SICI, SARD.
@@
  • oneone denotes one prosumer/producer
  • $\mathcal{P}$ is a set of prosumer/producer categories, where $\mathcal{P}\in\{\text{Public},\text{Residential},\text{SME},\text{NPO},\\ \text{Standalone producing installation}\}$.
@@
Similarly, the self-consumption level of 49.1% was identified as the average across all RECs for which such data were available. Consequently, our engineering model is based on a self-consumption range of 45%–50%–55%, where the central value corresponds to the real-world situation, the lower bound reflects a slightly more pessimistic scenario, and the upper bound represents a more optimistic scenario that reflects the potential deployment of battery energy storage systems (BESS). However, only 3.6% of the RECs currently report having plans to install BESS. Although BESS would enable both a higher self-consumption rate and increased remuneration for shared energy, long payback periods inhibit their adoption.
Another input parameter is the ’most common building type’, which is essential to select realistic prosumer load profiles in the engineering model. For example, public prosumers may be represented by the load profile of a school or a sports facility, while SME prosumers may be modelled using the load profiles of commercial buildings or hotels. Clearly, different load profiles yield different results. Therefore, identifying the most common building type is critical for producing outputs that closely reflect reality without introducing excessive complexity by modelling all possible building types. This parameter was derived from our database through manual counting of building types for different categories of prosumers.

### v2

@@ 2 Literature review
This Section reviews the relevant literature on the topic, by exploring two main areas [^10] We also discuss the studies that analyze the effect of prosumers on electricity systems in Section 5. However, since they do not constitute the central role to our paper, unlike studies that focus on the effect of REC deployment on the system, we omit reviewing them in the main text although including them in the Appendix 7.4.:
@@ 2.1 Impact of RECs on electricity systems
Most studies use a theoretical narrative-based approach to discuss the systemic impacts of RECs (Biggar and Hesamzadeh, 2022; Del Pizzo et al., 2022; Parag and Sovacool, 2016; Robinson and Del Guayo, 2022; Di Silvestre et al., 2021) and only few empirical economic studies exist (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025) due to the novelty and previously small scale of the phenomenon. Several studies indicate systemic benefits of REC deployment. Encouraging investment in embedded generation at a larger scale increases economic efficiency (Biggar and Hesamzadeh, 2022). Locating RECs in places with under-supply of electricity reduces transmission grid costs and losses (Robinson and Del Guayo, 2022). RECs reduce the need for an additional large-scale generation due to the merit-order effect, which in turn reduces the wholesale price and the need for additional transmission infrastructure (Robinson and Del Guayo, 2022; Backe et al., 2022). Robinson and Guayo (2022) argue that RECs can provide ancillary services if aggregated into virtual power plants, thereby replacing conventional sources which are more expensive to operate. When RECs are aggregated and utilize battery energy storage systems (BESS), they could reduce peak heat load and total heat demand in highly electrified systems (Backe et al., 2022). New ESCOs and aggregators could emerge as facilitators between RECs and the grid, therefore enhancing residential and commercial energy efficiency efforts (Parag and Sovacool, 2016). Finally, multiple social benefits (Biggar and Hesamzadeh, 2022; Robinson and Del Guayo, 2022) and improved competitive pressure on other market agents (Robinson and Del Guayo, 2022) are also important systemic benefits.
@@
At the same time, other studies point to systemic detriments of REC deployment. Biggar and Hesamzadeh (2022) and Robinson and Guayo (2022) argue that a REC may arbitrate [^11] Arbitrage effect emerges when collective generators and consumers/prosumers have distinct selling and buying prices. When they unite into an energy community, generator now can supply energy directly to members of the REC while latter remunerate the former with a payment that is higher than conventional price for selling into the grid but lower than a conventional buying price for consumers. As a result, both parties win while the regulatory effect of the price has vanished. The buying price for consumers from local generators can be implicitly lowered compared to buying price from a central market if the distribution/transmission charge is distance-based. between retail and selling prices in cases when it acts as a united entity which purchases part of its electricity externally from a single retailer. As a consequence, the effect of non-uniform tariffs (time-of-use and locational tariffs) is eliminated, which could make the system operation costlier and potentially may lead to the further expansion of the network. Biggar and Hesamzadeh (2022) indicate that retail customers do not typically face the correct incentives to use energy according to market and grid situations. When customers unite into RECs it may further amplify the impact of such inefficient tariffs. The same study (2022) argues that the reduction in network exploitation exacerbates network operator’s ‘death spiral’ effect [^12] Death spiral effect describes a situation when distributed energy resources lead to prosumers paying less with volumetric distribution charges. Consequently, system operators (DSOs/TSOs) become under-paid for its previous infrastructure investment, which is usually higher than a decreased usage of a network due to self-consumption. System operators shift these costs instead onto other consumers, which in turn stimulate them to transform into prosumers as well, thereby shrinking the customer base iteratively.. Robinson and Guayo (2022) and Parag and Sovacool (2016) say that locating RECs in places with over-supply of electricity increases transmission grid costs and losses due to increased electricity injection where congestion bottlenecks are present. In addition, small capacity renewable generators are costlier for the system because they are more expensive by LCOE than large capacity renewable generators (Robinson and Del Guayo, 2022). Parag and Sovacool (2016) point that RECs might erode sensitive protections on privacy. As we can see, the literature points to multiple potential harms to the system and stakeholders [^13] Since our study aims to analyze systemic impacts, we skip the literature review of impacts on the individual stakeholders in the main text. Instead, we include the review of these impacts in Appendix C. from the deployment of RECs. However, these potential harms relate to a variety of REC types and not all are necessarily applicable to the ‘community collective generation’ type deployed in Italy [^14] Koltunov (2025a) can be consulted for an in-depth discussion on the impacts of individual REC business models on the electricity system and its stakeholders..
@@
Fuentes Gonzalez et al. (2022) utilize a game theoretical approach with optimization modelling for a simplified three-nodal Chilean market. The study reveals that the nodal price decreases from 65.77 \$/MWh to 65.4 \$/MWh when RECs are deployed in a scenario with 9.24 MW of installed capacity with 150 MW peak nodal load. Moreover, the price drops further up to 64 \$/MWh if a REC’s installed capacity gets higher, though contingent upon the disposable income of REC members. The reduction in the nodal price suggests that the effect of RECs can also be positive on the non-member consumers. An increase in social welfare can be observed when RECs are involved in energy production (14,221 \$/hour with RECs versus 14,596 \$/hour without REC) due to the avoidance of high generation costs at nodes without REC, as well as the less frequent need for transmission expansion. This leads to a reduction in the nodal price that triggers an increase in the quantity demanded from 105.79 MWh to 115.03 MWh with an energy community. Quantities demanded in their static model are simulated for a single representative hour while prices are simply derived from linear demand functions. Consequently, this approach does not take into consideration a full interplay of technologies and variability over multiple time periods that are inherent in pay-as-clear zonal wholesale auctions of the real world.
@@
Sarfarazi et al. (2020) also use a game-theoretical approach with optimization and dynamic programming to simulate the interaction of a stylized German REC with its retailer and a wholesale market. In the first scenario, a community storage is owned by the retailer who pursues the profit maximization objective. Here, real-time buying and feed-in selling tariffs do not decrease the quantities dispatched at the supply side of the wholesale market by RECs compared to static buying tariffs. In the latter case, community storage decreases the energy dispatched at the supply side of the wholesale market. In general, the retailer’s profit-maximization scenario leads to welfare gains both for the system and a REC. Conversely, if a retailer pursues a REC self-sufficiency maximization objective and not a profit maximization objective, REC imports less electricity from the public grid, therefore, less of grid charges are collected for the DSOs and less of levies to the general budget. Sarfarazi et al. (2020) focus primarily on the ”REC-retailer” interactions on a micro scale and do not investigate REC deployment on a macro scale, merit-order effect (MOE) and complexities of a DA wholesale market.
@@
Boccard and Goetz (2025) find that the impact of RECs on the grid can be diverse. When a PV system is sized at 25% of the aggregated load of REC residential members the annual power exchanges with the grid drop from 70 MWh (no PV scenario) to 59 MWh (14% decrease) in Northern Spain. This is a significant finding, which indicates that a DSO could manage the deployment of RECs and generally prosumers without having to raise tariffs and compromising on grid reliability. When the PV capacity is increased to cover 50% of an aggregated load, the power exchanges with the grid decrease by 7%. Instead, in the scenario where a PV capacity covers 100% of the member load, power exchanges with a distribution grid raise up to 29%. The situation worsens in the northern irradiance region (Germany), whereas betters in the southern irradiance location (Portugal). In contrast, in the scenario with heterogeneous public prosumers (schools and hospitals) that install PV systems covering 50% and 100% of the load, the power exchanges reduce by 8% and 13%, respectively, compared to scenario without a PV. Instead, the 25% PV coverage of public loads rises power exchanges with the grid by 10%. The authors argue that the deployment of small-scale PV systems in residential RECs could decrease operational costs for the grid, while large-scale PVs force costly upgrades within residential buildings and across a distribution grid, and therefore better suit RECs composed of public prosumers.
@@
Last, Backe et al. (2022) utilized two large optimization models to explore the effect of the REC deployment on the additional capacity needs in a 2060 carbon-neutral Europe. The authors find that the REC deployment reduces both electricity and heating costs and lessens the need for national capacity expansion by 50-60 GW in six EU countries. However, they do not quantify the MOE or REC interaction with the DA market, openly stating that this is beyond the scope of their paper.
@@ 3 Methods and data
The first stage of our methodological approach begins with mapping all Italian RECs to derive input parameters for subsequent engineering modeling. The model’s output is used to project REC deployment for a year 2027 according to three scenarios, which provides the input variables for the second stage. In the second stage, we first simulate a market equilibrium with RECs in 2027. Then, we construct a counterfactual scenario of market equilibrium without RECs. As the most recent annual data on day-ahead market bids is available only for 2024, and the policy target is set for 2027, we assume that the wholesale market conditions in 2027 are identical to those of 2024 for simplification purposes. This assumption does not compromise our modeling objective, which is to introduce a new methodology for estimating the impact of REC deployment on wholesale market equilibrium. Our aim is not to forecast the actual impact of REC deployment in the policy-target year. Nevertheless, policy-relevant insights can be drawn from the modeled scenarios. In the first stage of methodology, we used a combination of software applications: MS Excel (mapping), Matlab (engineering modeling), R (projection). In the second stage, R was utilized as the main modeling software.
@@ 3.1.1 Comprehensive Mapping of Italian RECs
  • the data portal of GSE [^16] The state-owned renewable energy agency,
  • the annual reports of Legambiente [^17] A national environmental non-governmental organisation, i.e. NGO.,
@@
  • the data portal NeXt [^18] NeXt — the civil society network comprising the majority of Italian third-sector and public bodies working in the REC field ESG,
@@
  • the data portal Sinergie Condivise [^19] Developed by a banking foundation ‘Fondazione Compagnia di San Paolo‘ in collaboration with regional and municipal governments and universities in the Piedmont Region,
  • publicly available business plans of individual RECs,
  • academic publications,
  • websites of REC developers,
@@
  • websites of news agencies, regional/municipal authorities, and individual initiatives.
We refer to Italian electricity market zones using acronyms: North – NORD, Central North – CNORD, Central South – CSUD, Calabria – CALA, South – SUD, Sicily – SICI, Sardinia – SARD. The prosumer’s categories (public, residential, SMEs, NPOs) were determined based on subjects that are entitled to become a member of REC by Italian law (ARERA, 2022), namely public entities, private citizens, SMEs, and NPOs. Our final database contains 34 variables for 362 RECs in Italy, of which 184 are in the operational phase and 178 are in the design phase as of January 2025. However, for the purposes of this study, we used only 20 variables. Zhu et al. (2025) also built the database of new Italian RECs for the purpose of their study. In addition to the diverse scope of variables collected (due to different research objectives), another major difference lies in the development status of the RECs. Our database includes both operational RECs and those that were in the project design stage of development. Inclusion of RECs in the project design stage allowed us to have a more holistic perspective on future trends, which is discussed in further detail in Section 3.1.3. In turn, the database collected by GSE (2025b) though is complete but contains a very small set of variables  [^20] The recent map published by RSE (2025) is primarily based on the GSE data, although contain more variables than the original database, but still does not include critical variables (i.e. related to categories of members) used for our research.. However, neither of the other databases contain information about a category of members nor about the shares of installed PV capacity by prosumer categories. In contrast, our database contains these variables because it was collected specifically for the purpose of this study. A detailed comparison of three databases can be found in Table 3.1.1.
@@
Comparison between the databases Variablea Our database Zhu et al. (2025) GSE (2025b) Total number of RECs, from which: 362 212 344 Operational RECs 184 212 344 Design-phase RECs 178 0 0 Location 362 212 344 Population (of municipality) 85 Climatic zone 85 Market zone 362 Generator’s capacity 300 212 344 Type of technology installed/planned 323 85 Number of membersb 283 344 Category of members 219 Shares of installed PV capacity by category of prosumers 271 Technical indicatorsc 147 BESS availability 13 7 EVs and/or e-charging availability 38 3 Self-consumption level 82 Economic indicatorsd 105 Number of prosumer buildings by type 158 55 Number of consumer buildings not counted 41 Investment source 252 CAPEX costs 106 Legal form 180 344 REC builders and promoters 204 DSO operating a primary cabin 344 Last update January 2025 February 2025 October 2025 Data sources GSE, Legambiente, RSE, Sinergie, Condivise, NeXt ESG, business plans, academic publications, websites of builders, other websites GSE, Legambiente, RSE, academic publications own
@@
    Some reported variables represent a set of variables in actual database. For example, ”Shares of installed PV capacity by category of prosumers” contain five variables in the database, etc.
  • b
@@
  • $p$ denotes a category of prosumer/producer, where $\mathcal{P}\in\{\text{Public},\text{Residential},\text{SME},\text{NPO},\\ \text{Standalone producing installation}\}$.
  • $z$ denotes an electricity market zone, where $Z\in\{\text{NORD},\text{CNORD},\text{CSUD},\text{CALA},\\ \text{SUD},\text{SICI},\text{SARD}\}$.
  • oneone denotes one prosumer/producer.
@@
Similarly, the self-consumption level of 49.1% was identified as the average across all RECs. Consequently, our engineering model is based on a self-consumption range of 45%–50%–55%, where the central value corresponds to the real-world situation, the lower bound reflects a slightly more pessimistic scenario, and the upper bound represents a more optimistic scenario that reflects the potential deployment of BESS. However, only 3.6% of the RECs currently report having plans to install BESS. Although BESS would enable both a higher self-consumption rate and increased remuneration for shared energy, long payback periods inhibit their adoption.
Another input parameter is the ’most common building type’, which is essential to select realistic prosumer load profiles for the engineering model. For example, public prosumers may be represented by the load profile of a school or a sports facility, while SME prosumers may be modelled using the load profiles of commercial buildings or hotels. Clearly, different load profiles yield different results. Therefore, identifying the most common building type is critical for producing outputs that closely reflect reality without introducing excessive complexity by modelling all possible building types. This parameter was derived from our database through manual counting of building types for different categories of prosumers.